Estelle deposits R1000 into a savings account paying 6.5% interest per year compounded quarterly, after three and a half years she withdraws R 500 from hte account and deposits it into another account paying 12% interest per year compounded semi annually, how much is the total amount accrued from both accounts two years after making the second deposit?
Answer by Boreal(15235)

The first 3.5 years
P=Po{1+r/n}^nt
=1000(1+.065/4)^14
=1253.16, rounding at end.
withdraw 500 and now has initial of 753.16 for 2 years, so 753.16(1+(0.065/4))^8=856.82 account 1
account 2 is 500(1+(.12/2)^4=631.24
Total is R1488.06